Write down formula for maximum height of projectile motion. And relation to maximum height and range.
1
Expert's answer
2020-11-02T09:26:33-0500

Solution

In case of maximum height projection angle is 90°

So Maximum height can be expressed by following formula

"H_{max}=\frac {v^2}{2g}" ....................... Eq 1

Where v is projection velocity.

Now for maximum range angle must be 45° so

Maximum

"R_{max}=\frac{v^2}{g}" ........................... eq 2

Using equation 1 and 2

Now relationship between maximum height and range can be written as


"H_{max}=\frac {R_{max}}{2}"